OpenStudy (anonymous):

what is (5x^5 - 5x^3 - 10x^2 - 80x) divided by(x + 2)=?

OpenStudy (anonymous):

let p(x) = 5x^5 - 5x^3 - 10x^2 - 80x p(-2) = -160 + 40 -40 + 160 = 0 so x+2 is a factor

OpenStudy (anonymous):

\[\Large \frac{5x^5-5x^3-10x^2-80x}{x+2}=\frac{5x(x+2)(x^3-2x^2+3x-8)}{x+2 }...\] .....

OpenStudy (anonymous):

calculator ... lol

OpenStudy (anonymous):

your calculator can do that? awesome

OpenStudy (anonymous):